# Inflection Point Inflation in Supergravity

### v1

@@ 3.1 Analytic Analysis of the Model
We also note that the curvature of the potential is negative for an extended range of $\phi$ below $\phi_{0}$. In the non-SUSY case, this holds for $\phi/\phi_{0}\in[1/3,1]$, independently of the value of $\phi_{0}$; in the SUGRA case with $\phi_{0}\ll 1$, this region extends to $\phi/\phi_{0}\in[1/4,1]$. The minimum of the curvature occurs at $\phi\simeq 0.54\phi_{0}$, closer to the origin than in the non-SUSY case, with a value just below $-0.23~{}m_{\phi}^{2}$, smaller in magnitude than in the non-SUSY case. The latter reduces the tachyonic instability while the former increases it. Hence, we expect the non-perturbative effects after inflation studied in Drees:2025iue for the non-SUSY case would be similar in the SUGRA version of the model.
@@ 3.2 Numerical Results of the Model
[…]
The running of spectral index $\alpha$ is shown in the bottom right frame. It remains independent of $\phi_{0}$ even for $\phi_{0}\geq 1$. In contrast to the non-SUSY version of the model, $\alpha$ strongly depends on $N_{\textrm{cmb}}$. When $N_{\textrm{cmb}}=45$ $\alpha\approx-0.0032$. Increasing $N_{\textrm{cmb}}$ reduces the absolute value of $\alpha$, reaching $\alpha\approx-0.0013$ for $N_{\textrm{cmb}}=65$.
Finally, the bottom left frame of Fig. 2 shows the relationship between the tensor-to-scalar ratio $r$ and $\phi_{0}$. It follows the same pattern as the Hubble scale. However, since we find $r<10^{-7}$, a positive detection of tensor modes by current or near-future experiments would exclude our model.
@@ 3.3 SUSY Breaking by a Polonyi Field
$\begin{split}W&=B\Phi^{2}+C\Phi^{3}+D\Phi^{4}+\mu M_{\textrm{pl}}(Z+\beta_{\rm P})\\ &=W_{\textrm{I}}+W_{\textrm{P}}\end{split}$ (32)
where $\mu$ essentially sets the SUSY breaking scale; we explicitly include a factor $M_{\textrm{pl}}$ here to ensure the correct dimension of $\mu$. Both $\Phi$ and $Z$ are complex fields. As before, we want the real part of $\Phi$ to be the inflaton field $\phi$. $Z$ is the Polonyi field whose vacuum expectation value $\langle Z\rangle$ is the only source of SUSY breaking after inflation, when $\langle\phi\rangle=0$. The SUSY breaking with vanishing vacuum energy requires $\beta_{\rm P}=2-\sqrt{3}$, which gives the gravitino mass $m_{3/2}=\mu e^{2-\sqrt{3}}$ when $Z$ stays at the SUSY breaking minimum at $Z=\langle Z\rangle=(\sqrt{3}-1)M_{\textrm{pl}}$.
@@
We require that the Polonyi field does not change the inflation potential significantly, which means $|W_{\textrm{I}}|\gg|W_{\textrm{P}}|$. The existence of the SUSY breaking term will not alter the slow roll parameters significantly if
@@
$V_{\mu}(\phi)=\frac{{\rm e}^{\frac{1}{2}\phi^{2}}}{4}\mu\left((-4\beta_{\rm P}B\phi^{2}+2\beta_{\rm P}(B+D)\phi^{4}+\sqrt{2}\beta_{\rm P}C\phi^{5}+\beta_{\rm P}D\phi^{6})+\mu(4-12\beta_{\rm P}^{2}+2\beta_{\rm P}^{2}\phi^{2})\right)\,.$ (35)
@@
$\begin{split}\epsilon_{\mu}\approx&\sqrt{2\epsilon_{\rm cmb}}\frac{\left(8\sqrt{3}-13\right)\tilde{\mu}^{2}\phi_{0}+\left(4\sqrt{3}-8\right)\tilde{\mu}\phi_{0}^{3}}{2\phi_{0}^{6}}\,,\\ \eta_{\mu}\approx&\frac{(8\sqrt{3}-13)\tilde{\mu}^{2}+(5\sqrt{3}-10)\tilde{\mu}\phi_{0}^{4}}{2\phi_{0}^{6}}\,,\end{split}$ (36)
@@
On the other hand, if $W_{\textrm{P}}\gg W_{\textrm{I}}$, the Polonyi field will stay at $Z=\sqrt{3}-1$. In this case, the inflection point conditions $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$ up to $\phi_{0}^{2}$ read:
$\begin{split}V^{\prime}&=0\Rightarrow 8\tilde{B}^{2}+18\sqrt{2}\tilde{B}\tilde{C}\phi_{0}-4\left(\sqrt{3}-2\right)\tilde{B}\tilde{\mu}+18\tilde{C}^{2}\phi_{0}^{2}-3\sqrt{2}\left(\sqrt{3}-3\right)\tilde{C}\tilde{\mu}\phi_{0}+2\tilde{\mu}^{2}=0\,,\\ V^{\prime\prime}&=0\Rightarrow 4\tilde{B}^{2}+18\sqrt{2}\tilde{B}\tilde{C}\phi_{0}-2\left(\sqrt{3}-2\right)\tilde{B}\tilde{\mu}+27\tilde{C}^{2}\phi_{0}^{2}-3\sqrt{2}\left(\sqrt{3}-3\right)\tilde{C}\tilde{\mu}\phi_{0}+\tilde{\mu}^{2}=0\,,\end{split}$ (38)
with the following solutions:
$\tilde{B}\approx 0.4952\tilde{\mu}\,,\quad\tilde{C}\approx-\frac{0.4996\tilde{\mu}}{\phi_{0}}\,.$ (39)
The potential at the inflection point is then:
@@
This expansion is similar in structure from the previous cases. Following the same procedure and using $n_{s}=0.9659,\,N_{\textrm{cmb}}=65$, we have:
@@
If the Polonyi field already sits in its SUSY breaking minimum during inflation, all relevant energy scales, i.e., Hubble scale $H_{\inf}$, the SUSY breaking scale $\mu$, and inflaton mass $m_{\phi}$, are completely determined by the position of inflection point $\phi_{0}$. The scaling of $H_{\inf}$ and $m_{\phi}$ with $\phi_{0}$ is also as in the non-SUSY version of the model, or as in the SUGRA model without Polonyi sector. The new feature is that $\phi_{0}$, or $H_{\textrm{inf}}$, also determines $\mu$; again demanding $\mu>1$ TeV therefore implies $\phi_{0}>2\times 10^{-4}$ in this set-up. This strong correlation can only be relaxed by lifting the Polonyi field away from the SUSY breaking point $\sqrt{3}-1$.
In Fig. 4, we show how different quantities depend on $\phi_{0}$. For $\phi_{0}\ll 1$ this is described by eqs.(43) and (42), while $\phi_{0}\geq 1$ can again only be treated numerically. As before, we fix the spectral index $n_{s}$ and the number of e-folds $N_{\textrm{cmb}}$, and additionally $\tilde{\mu}$ for better illustration.
@@
So far we fixed $\tilde{\mu}=0.01$. For $\phi_{0}\ll 1$ this choice is in fact irrelevant, since the physical parameters $B,\,\delta B,\,C$ and $\mu$ are all fixed uniquely for given $\phi_{0}$, see eqs.(39), (42) and (43). However, we see in Fig. 5, which shows the relation between $\mu$ and $H_{\inf}$, that this is no longer true for $\phi_{0}\geq 0.1$. Nevertheless, for a given Hubble scale during inflation there will be a maximum SUSY breaking scale it can host, corresponding to the case where SUSY is already broken by the Polonyi field:
$\mu<\frac{4}{\phi_{0}}H_{\inf}\approx 3\times 10^{4}\left(\frac{H_{\inf}}{\textrm{GeV}}\right)^{2/3}\textrm{GeV}\,.$ (44)
Once the Polonyi field moves away from the SUSY breaking point, one will have more freedom to set the SUSY breaking value, depending on the position of the Polonyi field during inflation. If the Polonyi field stays at the origin and only perturbs the potential, the SUSY breaking scale would simply be $\mu=\tilde{\mu}\times D$, where $D$ can be treated as a constant. This explains why the relative ratio of three different cases when they deviate from the straight line in Fig. 5 is almost a constant.
@@
Figure 5: The scale of SUSY breaking $\mu$ vs. the inflationary Hubble scale $H_{\inf}$ on a log-log scale. Different colors represent different choices of relative scale $\tilde{\mu}$. The straight line is the Polonyi field dominated case, where the SUSY breaking scale only depends on the inflection point positions. The right, flat region is an inflation field-dominated region, where the SUSY breaking scale depends linearly on the relative scale $\tilde{\mu}$.
@@
4 Summary
In this paper we revisited the renormalizable inflection point inflation model in the SUGRA framework. We adopted the minimal assumption that only one canonical field drives inflation. While supersymmetry protects the flatness of the potential from radiative corrections, local SUSY or SUGRA also modifies the potential through non-renormalizable terms. These new terms contribute to slow roll parameters on an equal footing. As in the non-supersymmetric case the shape of the potential is determined by the position $\phi_{0}$ of the inflection point, which is a free parameter of our model. When fixing the well-constrained power spectrum of curvature perturbations and its spectral index, we find $\phi_{0}$ controls the tensor-to-scalar ratio $r$, the running of the spectral index $\alpha$, the Hubble scale during inflation $H_{\textrm{inf}}$, and the physical inflaton mass $m_{\phi}$.
For $\phi_{0}\ll 1$ a perturbative treatment is possible. In this case, $r,\,H_{\textrm{inf}}$ and $m_{\phi}$ are monomial functions of $\phi_{0}$ and reach their maximum around $\phi_{0}\approx 1$. The running of the spectral index $\alpha$ is almost independent of $\phi_{0}$ but depends more strongly on $N_{\textrm{cmb}}$ than in the non-supersymmetric, renormalizable version of the model. The tensor-to-scalar ratio $r$ is always smaller than $10^{-7}$, which is below the sensitivity of any current or planned experiments. The running of the spectral index, which lies in order of $\mathcal{O}(-10^{-3})$, might be probed by the next generation of CMB experiments. The predictions of this SUGRA model are quite similar to those from the renormalizable model. All observables have the same scaling with respect to $\phi_{0}$. Thus, even though the SUGRA potential contains terms up to $\phi^{6}$ while the renormalizable potential only has terms up to $\phi^{4}$, it still provides a relatively reliable estimate of inflationary quantities.
The main difference between the SUGRA case and the renormalizable case appears when $\phi_{0}$ exceeds $1$. In this region the exponential factor $e^{1/2\phi^{2}}$ in the SUGRA case becomes large, which suppresses $r,\,H_{\textrm{inf}}$ and $m_{\phi}$. The energy scales are therefore bounded from above: $H_{\textrm{inf}}<10^{11}\ \textrm{GeV}$ and $m_{\phi}<10^{12}\ \textrm{GeV}$. The renormalizable potential is not able to capture this behavior, which leads to a very different prediction of inflationary observables in this large field scenario.
We further added a SUSY breaking Polonyi sector to the model. If the SUSY breaking scale is much smaller than the Hubble scale, the Polonyi field will stay at the origin and serve as a perturbation to the field. When these two energy scales become comparable, the Polonyi field will move away from the origin and modify the inflation potential. These effects lead to a nontrivial bound between the SUSY breaking scale and the inflation scale. We find that for a TeV scale SUSY breaking, we need the Hubble scale to be larger than $1$ GeV.
It has been pointed out that in the KKLT model, the Hubble scale is always smaller than the gravitino mass $m_{3/2}$ (or SUSY breaking scale $\mu$Kallosh:2004yh . In our model, we find a slightly different conclusion: in some regions, the Hubble scale can be larger than the gravitino mass. In such a scenario supersymmetry will protect the inflaton potential from loop corrections. This should allow larger couplings of the inflaton to Standard Model (super)fields, and thus larger reheat temperature than in the non–supersymmetric version.

### v2

@@ 3.1 Analytic Analysis of the Model
We also note that the curvature of the potential is negative for an extended range of $\phi$ below $\phi_{0}$. In the non-SUSY case, this holds for $\phi/\phi_{0}\in[1/3,1]$, independently of the value of $\phi_{0}$; in the SUGRA case with $\phi_{0}\ll 1$, this region extends to $\phi/\phi_{0}\in[1/4,1]$. The minimum of the curvature occurs at $\phi\simeq 0.54\phi_{0}$, closer to the origin than in the non-SUSY case, with a value just below $0.23~m_{\phi}^{2}$, smaller in magnitude than in the non-SUSY case. The latter reduces the tachyonic instability while the former increases it. Hence, we expect the non-perturbative effects after inflation studied in Drees:2025iue for the non-SUSY case would be similar in the SUGRA version of the model.
@@ 3.2 Numerical Results of the Model
Figure 1: Rescaled inflation potential for different choices of the location of the inflection point $\phi_{0}$. Here $V_{0}=V(\phi_{0})$ is the value of the potential at the inflection point. Blue, orange, green, and red curves corresponding to $\phi_{0}=0.1,\,1,\,3$ and $5$, respectively.
@@
The relation between inflaton mass $m_{\phi}$ and inflection point $\phi_{0}$, shown in the top right frame, follows a similar pattern as the Hubble scale when $\phi_{0}$ is small, albeit with different power. However, the inflaton mass drops dramatically for $\phi_{0}>1$, due to the exponential suppression discussed at the end of the previous section. When $\phi_{0}\approx 10$, the inflaton mass could be as low as $1\ \textrm{GeV}$. This tiny value differs from the Hubble scale by more than nine orders of magnitude. Thus our model offers a way to separate the Hubble and inflaton mass scales.
The running of the spectral index $\alpha$ is shown in the bottom right frame. It remains independent of $\phi_{0}$ even for $\phi_{0}\geq 1$. In contrast to the non-SUSY version of the model, $\alpha$ strongly depends on $N_{\textrm{cmb}}$. When $N_{\textrm{cmb}}=45$, $\alpha\approx-0.0032$. Increasing $N_{\textrm{cmb}}$ reduces the absolute value of $\alpha$, reaching $\alpha\approx-0.0013$ for $N_{\textrm{cmb}}=65$.
Finally, the bottom left frame of Fig. 2 shows the relationship between the tensor-to-scalar ratio $r$ and $\phi_{0}$. It follows a similar pattern as the Hubble scale. However, since we find $r<10^{-7}$, a positive detection of tensor modes by current or near-future experiments would exclude our model.
In principle one can thus distinguish the small and large field scenarios through the tensor-to-scalar ratio. If we restrict the inflection point position so that it should not be much larger than the Planck scale, $\phi_{0}<10$, then a small tensor to scalar ratio $r<10^{-10}$ would favor the small field scenario. If the tensor-to-scalar ratio is between $10^{-8}$ and $10^{-10}$, we need to check further the inflation mass. It cannot be measured directly; however, generally speaking, a large reheating temperature requires a large inflation mass, which can only be realized in the small field version of our model. The reheating dynamics can in principle be probed via the observation of primordial high-frequency gravitational waves Bernal:2023wus ; Xu:2024fjl ; Jiang:2024akb ; Xu:2025wjq . Unfortunately none of these observations are feasible with near–future technology.
@@ 3.3 SUSY Breaking by a Polonyi Field
[…]
where $\mu$ essentially sets the SUSY breaking scale; we explicitly include a factor $M_{\textrm{pl}}$ here to ensure the correct dimension of $\mu$. Both $\Phi$ and $Z$ are complex fields. As before, we want the real part of $\Phi$ to be the inflaton field $\phi$. $Z$ is the Polonyi field whose vacuum expectation value $\langle Z\rangle$ is the only source of SUSY breaking after inflation, when $\langle\phi\rangle=0$. SUSY breaking with vanishing vacuum energy requires $\beta_{\rm P}=2-\sqrt{3}$, which gives the gravitino mass $m_{3/2}=\mu e^{2-\sqrt{3}}$ when $Z$ stays at the SUSY breaking minimum at $Z=\langle Z\rangle=(\sqrt{3}-1)M_{\textrm{pl}}$.
@@
We require that the Polonyi field does not change the inflation potential significantly, which means $|W_{\textrm{I}}|\gg|W_{\textrm{P}}|$ during inflation. The existence of the SUSY breaking term will not alter the slow roll parameters significantly if
@@
$V_{\mu}(\phi)=\frac{{\rm e}^{\frac{1}{2}\phi^{2}}}{4}\mu\left[(-4\beta_{\rm P}B\phi^{2}+2\beta_{\rm P}(B+D)\phi^{4}+\sqrt{2}\beta_{\rm P}C\phi^{5}+\beta_{\rm P}D\phi^{6})+\mu(4-12\beta_{\rm P}^{2}+2\beta_{\rm P}^{2}\phi^{2})\right]\,.$ (38)
@@
[…]
@@
On the other hand, if $|W_{\textrm{P}}|\gg|W_{\textrm{I}}|$, the Polonyi field will stay at $Z=\sqrt{3}-1$; this requires $|\mu|\gg|D|\phi_{0}^{4}$. Focusing on small field solutions, and imposing the slightly stronger condition $|\mu|\gg|D|\phi_{0}^{2}$, the inflection point conditions $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$ up to terms ${\cal O}(\phi_{0}^{2})$ read:
[…]
with the following solutions: [^8] This solution implies $B^{2}\ll C^{2}$ as long as $\phi_{0}\ll 1$, which a posteriori justifies dropping the quartic term $\propto B^{2}$ in eq.(13). Moreover, eqs.(42) and our assumption $|\tilde{\mu}|\gg\phi_{0}^{2}$ imply that the third quartic term in eq.(13), $\propto BD$, is also negligible. We note that without this additional assumption, higher than quartic powers of $\phi$ need to be considered, in which case no simple analytical treatment is possible.
$\tilde{B}\approx 0.4952\tilde{\mu}\,,\quad\tilde{C}\approx-\frac{0.4996\tilde{\mu}}{\phi_{0}}\,$, (42)
i.e. the ratio $B/C$ again needs to be tuned for inflection point inflation to occur. The potential at the inflection point is then:
@@
This expansion is similar in structure to the previous cases. Following the same procedure and using $n_{s}=0.9659,\,N_{\textrm{cmb}}=65$, we have:
@@
If the Polonyi field already sits in its SUSY breaking minimum during inflation, all relevant energy scales, i.e., Hubble scale $H_{\inf}$, the SUSY breaking scale $\mu$, and inflaton mass $m_{\phi}$, are completely determined by the position of inflection point $\phi_{0}$. The scaling of $H_{\inf}$ and $m_{\phi}$ with $\phi_{0}$ is again as in the non-SUSY version of the model, or as in the SUGRA model without Polonyi sector. The new feature is that $\phi_{0}$, or $H_{\textrm{inf}}$, also determines $\mu$; again demanding $\mu>1$ TeV therefore implies $\phi_{0}>2\times 10^{-4}$ in this set-up. This strong correlation can only be relaxed by lifting the Polonyi field away from the SUSY breaking point $\sqrt{3}-1$.
@@
In Fig. 4, we show how different quantities depend on $\phi_{0}$. For $\phi_{0}\ll 1$ and $\phi_{0}^{2}\ll\tilde{\mu}$ this is described by eqs.(46) and (45). For our choice $\tilde{\mu}=0.01$ the latter condition is stronger; indeed, we observe some deviations from eqs.(46) once $\phi_{0}\gtrsim 0.1$. As before, we fix the spectral index $n_{s}$ and the number of e-folds $N_{\textrm{cmb}}$.
@@
Figure 5: The scale of SUSY breaking $\mu$ vs. the inflationary Hubble scale $H_{\inf}$ on a log-log scale. Different colors represent different choices of relative scale $\tilde{\mu}$. The straight line is the Polonyi field dominated case, where the SUSY breaking scale only depends on the inflection point positions. The right, flat region is an inflaton field dominated region, where the SUSY breaking scale depends linearly on the relative scale $\tilde{\mu}$.
So far we fixed $\tilde{\mu}=0.01$. For $\phi_{0}^{2}\ll\tilde{\mu}$ this choice is in fact irrelevant, since the physical parameters $B,\,\delta B,\,C$ and $\mu$ are all fixed uniquely for given $\phi_{0}$, see eqs.(42), (45) and (46). However, we see in Fig. 5, which shows the relation between $\mu$ and $H_{\inf}$, that this is no longer true for $\phi_{0}\geq\sqrt{\tilde{\mu}}$. Nevertheless, this figure shows that for a given Hubble scale during inflation there will be a maximum SUSY breaking scale it can host, given by the analytical solution of eq.(46):
$\mu<\frac{10.5}{\phi_{0}}H_{\inf}\approx 2.3\times 10^{4}\left(\frac{H_{\inf}}{\textrm{GeV}}\right)^{2/3}\textrm{GeV}\,.$ (47)
Recall that this corresponds to a scenario where the Polonyi field remains at its current value throughout inflation. If it moves away from this value during inflation, one will have more freedom to set the value of $\mu$, depending on the position of the Polonyi field during inflation. In the opposite extreme, where the Polonyi field stays at the origin during inflation and only perturbs the inflaton potential, the SUSY breaking scale would simply be $\mu=\tilde{\mu}\times D$, where $D$ can be treated as a constant. This explains why the relative ratio of three different cases when they deviate from the straight line in Fig. 5 is almost a constant.
We conclude that only the light cyan region below the topmost line in Fig. 5 is accessible in our model. Different choices of $\tilde{\mu}$ will leave the straight line at different Hubble scales, and can thus populate this region, always keeping in mind that $\mu>1$ TeV is needed for phenomenological reasons. The resulting lower bound on the Hubble scale during inflation is around $1~\textrm{GeV}$. This bound is much smaller than the naive estimate of $10^{8}$ GeV we derived below eq.(40) from the requirement that the Polonyi sector can be treated as a small perturbation.
4 Summary
In this paper we revisited the inflection point inflation model in the SUGRA framework. We adopted the minimal assumption that only one canonical field drives inflation. While supersymmetry protects the flatness of the potential from radiative corrections, local SUSY or SUGRA also modifies the potential through non-renormalizable terms. These new terms contribute to slow roll parameters on an equal footing. As in the non-supersymmetric case the shape of the potential is determined by the position $\phi_{0}$ of the inflection point, which is a free parameter of our model. When fixing the well-constrained power spectrum of curvature perturbations and its spectral index, we find $\phi_{0}$ controls the tensor-to-scalar ratio $r$, the Hubble scale during inflation $H_{\textrm{inf}}$, and the physical inflaton mass $m_{\phi}$.
For $\phi_{0}\ll 1$ a perturbative treatment is possible. In this case, $r,\,H_{\textrm{inf}}$ and $m_{\phi}$ are monomial functions of $\phi_{0}$ and reach their maximum around $\phi_{0}\approx 1$. The running of the spectral index $\alpha$ is almost independent of $\phi_{0}$ but depends more strongly on $N_{\textrm{cmb}}$ than in the non-supersymmetric version of the model. The tensor-to-scalar ratio $r$ is always smaller than $10^{-7}$, which is below the sensitivity of any current or planned experiments. The running of the spectral index, which is $\mathcal{O}(-10^{-3})$, might be probed by the next generation of CMB experiments. The predictions of this SUGRA model are quite similar to those from the non-supersymmetric model. All observables have the same scaling with respect to $\phi_{0}$. Thus, even though the SUGRA potential contains terms up to $\phi^{6}$ while the non-supersymmetric potential only has terms up to $\phi^{4}$, the latter still provides a relatively reliable estimate of inflationary quantities.
The main difference between the SUGRA case and the non-supersymmetric case appears when $\phi_{0}$ exceeds $1$. In this region the exponential factor $e^{\phi^{2}/2}$ in the SUGRA case becomes large, which suppresses $r,\,H_{\textrm{inf}}$ and $m_{\phi}$. The energy scales are therefore bounded from above: $H_{\textrm{inf}}<10^{11}\ \textrm{GeV}$ and $m_{\phi}<10^{12}\ \textrm{GeV}$. The non-supersymmetric potential is not able to capture this behavior, which leads to a very different prediction of inflationary observables in this large field scenario.
We further added a SUSY breaking Polonyi sector to the model. If the SUSY breaking scale is much smaller than the Hubble scale, the Polonyi field will stay at the origin and serve as a perturbation to the field. When these two energy scales become comparable, the Polonyi field will move away from the origin and modify the inflation potential. In the opposite extreme the Polonyi field stays close to its post-inflationary value throughout. These effects lead to a nontrivial bound between the SUSY breaking scale and the inflation scale. We find that for a TeV scale SUSY breaking, we need the Hubble scale to be larger than $1$ GeV.